# Rethinking the Relationship between the Power Law and Hierarchical Structures

### v2

@@ Appendix C Alternative Estimators
To further support the reliability of our estimates, we compute $I_{\text{POS}}$, $I_{\text{tag}}$, and $J$ using six alternative estimators proposed by Chao and Shen (2003) (CS), Horvitz and Thompson (1952) (HT), Miller (1955) and Madow (1948) (MM), Nemenman et al. (2001) (NSB), Wolpert and Wolf (1995) (WW) and Zahl (1977) (ZA). For these computations, we use the Python package entropy-estimation (Arora et al., 2022).
@@
We also show the estimates of the CFIB $J$ obtained with the MM, NSB, and ZA estimators in Figures 13 (c), (d), and (e), respectively. The MM and ZA estimates continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, similar to the GR estimates in the main text. This means that these two estimates properly converge to zero, the true value of the CFIB in PCFGs. The convergence of the NSB estimates is rather unclear. We do not present results with CS, HT, and WW, because most of their estimates are negative and cannot be plotted on a log scale.
@@
Appendix D Context-free Independence Breaking

### v3

@@ Appendix D Alternative Estimators
To further support the reliability of our estimates, we compute $I_{\text{POS}}$, $I_{\text{tag}}$, and $J$ using six alternative estimators proposed by Chao and Shen (2003) (CS), Horvitz and Thompson (1952) (HT), Miller (1955) and Madow (1948) (MM), Nemenman et al. (2001) (NSB), Wolpert and Wolf (1995) (WW) and Zahl (1977) (ZA). For these computations, we use the Python code entropy-estimation implemented by Arora et al. (2022).
@@
We also show the estimates of the CFIB $J$ for the PCFG obtained with the MM, NSB, and ZA estimators in Figures 13 (c), (d), and (e), respectively. The MM and ZA estimates continue to decrease in the log-scale plot as $N_{\text{data}}$ increases, similar to the GR estimates in the main text. This means that these two estimates properly converge to zero, the true value of the CFIB in PCFGs. The convergence of the NSB estimates is rather unclear. We do not present results with CS, HT, and WW, because most of their estimates are negative and cannot be plotted on a log scale.
@@
Appendix E Context-free Independence Breaking